if the perimeter of a rectangular lawn is 116 m and the width is 6 m less than the length, then what are the lengths and width?
Answer by stanbon(75887)

if the perimeter of a rectangular lawn is 116 m and the width is 6 m less than the length, then what are the lengths and width?
-----------------------------
P = 2(length + width)
---
116 = 2(L + L-6)
58 = 2L-6
2L = 64
Length = 32 m
----
Width = 32-6 = 26 m
=================================
Cheers,
Stan H.
=============